# Geometric Bounds for Persistence

### v1

@@ 3.3 Bounds for Vietoris–Rips Lifespans via Urysohn Width
$d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})),$
where $\omega_{r}$ maps to $\omega$ by the homology level map induced by the inclusion, and $\kappa:\mathcal{X}\to L^{\infty}(\mathcal{X})$ is the Kuratowski embedding. Recall also (2.34) that
$\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))=\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))).$
To estimate VR-lifespans, we need to know how to upper-bound quantities like $\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$, for which Theorem 3.3 is not applicable: the set $\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$ is not compact (even after taking the closure). One way to deal with non-compactness is to intersect $\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$ with the closure of the convex hull of $\mathcal{X}$, like we did in 3.11. That theorem, together with the inequality $\operatorname{TW}_{k}(\cdot)\leq\operatorname{AW}_{k-1}(\cdot)$, will give us a bound on $\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$ in terms of the Alexandrov width of $\overline{\mathcal{N}}_{r}(\kappa(\mathcal{X})\subset\overline{\mathrm{conv}}(\kappa(\mathcal{X})))$. It turns out the latter can be interpreted in terms of the Urysohn width.
@@
In the inequality $\rho(\omega;\mathcal{X})\leq\frac{1}{2}\operatorname{UW}_{k-1}(\mathcal{X})$, the left-hand side is defined via the Kuratowski embedding in $L^{\infty}(\mathcal{X})$, and it is instructive to interpret the right-hand side as well, in a way compatible with the picture of $\kappa(\mathcal{X})$ sitting in $L^{\infty}(\mathcal{X})$. This interpretation goes back to V. Tikhomirov [Tik76]. Some references in English are [Gro83, Appendix 1, Proposition (D)] or [Bal21, Theorem 2.1.9 and Definition 2.3.1]; we provide here a short summary. The width $\operatorname{UW}_{k-1}(\mathcal{X})$ can be equivalently defined as the infimal number $\delta>0$ such that there is map $f:\mathcal{X}\to L^{\infty}(\mathcal{X})$ whose image is at most $(k-1)$-dimensional complex in $L^{\infty}(\mathcal{X})$, and such that $\|\kappa(x)-f(x)\|_{\infty}\leq\delta/2$ for all $x\in\mathcal{X}$. In a nutshell, Theorem 3.3 says that one can try killing higher homology of $\mathcal{X}$ by deforming $\kappa(\mathcal{X})$ to a low-dimensional complex inside $L^{\infty}(\mathcal{X})$, and if every point moves by some controlled distance, then we get an estimate for the filling radius. This interpretation can be shortly reformulated as follows:
@@
$\rho(\omega;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))\leq\frac{1}{2}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\kappa(\mathcal{X})\subset\overline{\mathrm{conv}}(\kappa(\mathcal{X})))).$
We will go further and estimate $\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$ even better, by replacing the convex hull with the tight span (2.27). The following result subsumes Theorem 3.3 when $r=0$. The specific realization of $\operatorname{E}(\mathcal{X})$ in $L^{\infty}(\mathcal{X})$ that was described in Section 2.2 has an isometric copy of $\mathcal{X}$ inside, and the corresponding embedding of $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$ agrees with the Kuratowski embedding. To simplify notation a bit, we omit $\kappa$ and simply write $\mathcal{X}\subset\operatorname{E}(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ in the rest of this section. Notice that the set $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$ is compact, as a closed subset of the compact metric space $\operatorname{E}(\mathcal{X})$.
@@
For any compact metric space $\mathcal{X}$, any integer $k\geq 1$, any real number $r\geq 0$, and any homology class $\omega\in\mathrm{H}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))$,
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))\leq\frac{1}{2}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
@@
We have a chain of distance-preserving embeddings $\mathcal{X}\subset\operatorname{E}(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ as well as $\mathcal{X}\subset\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. All distances below are measured using the $L^{\infty}$-norm. We need to bound $\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))$.
Recall the following property of the tight span: there exists a $1$-Lipschitz retraction $\pi:L^{\infty}(\mathcal{X})\to\operatorname{E}(\mathcal{X})$. Our first observation is that $\pi$ maps $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ to $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$. Indeed, take any point $y\in\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Since $\|y-x\|<r$ for some $x\in\mathcal{X}$, and $\pi$ is $1$-Lipschitz retraction, it follows that $\|\pi(y)-x\|=\|\pi(y)-\pi(x)\|\leq\|y-x\|<r$. Therefore, $\pi(\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))=\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$. Denote the reverse inclusion $\iota:\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\to\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Since both $y$ and $\pi(y)$ lie in the open radius $r$ ball centered at $x$, the whole straight line segment between them lies in that ball and in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Therefore, a cycle representing $\omega$ can be continuously deformed to a cycle representing $\iota_{*}\pi_{*}\omega$ by letting each $y$ in the support of $\omega$ slide along the straight line segment towards $\pi(y)$. This homotopy happens entirely in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$, and therefore, $\omega=\iota_{*}\pi_{*}\omega$.
@@
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))\leq\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})).$
Indeed, any cycle $S$ representing $\pi_{*}\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))$ lies entirely in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ and represents there the homology class $\iota_{*}\pi_{*}\omega=\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))$. If $S$ bounds in the neighborhood of $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$ of certain radius, then it bounds even in a smaller neighborhood of $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$.
Using the evident $1$-Lipschitz restriction map $L^{\infty}(\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))\to L^{\infty}(\mathcal{X})$, it is easy to show that
$\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))\leq\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
(In fact, the equality holds here, but we only need the inequality.) To show this inequality, we can just use that $1$-Lipschitz map to send cycles and filling chains in $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))$ to cycles and filling chains in $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})$.
The next step follows easily from the definition of the filling radius:
$\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))=\rho(\pi_{*}\omega;\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
@@
$\rho(\pi_{*}\omega;\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))\leq\frac{1}{2}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
@@
$d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})),$
@@
$\lim_{r\to b_{\omega}+0}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))=\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$

### v2

@@ 3.3 Bounds for Vietoris–Rips Lifespans via Urysohn Width
$d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho\big{(}\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)},$
where $\omega_{r}$ maps to $\omega$ by the homology level map induced by the inclusion, and $\kappa:\mathcal{X}\to L^{\infty}(\mathcal{X})$ is the Kuratowski embedding. Recall also (2.38) that the right-hand side can be rewritten a bit shorter:
$\rho\big{(}\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)}=\rho\big{(}\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\big{)}.$
This follows from the following lemma.
Lemma 3.14.
Let $\mathcal{Z}$ be a subspace of $L^{\infty}(\mathcal{X})$ containing the Kuratowski image $\kappa(\mathcal{X})$. Then, for any homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{Z})$, the following is true:
$\rho(\omega;\mathcal{Z}\subset L^{\infty}(\mathcal{X}))=\rho(\omega;\mathcal{Z}).$
Proof.
The inequality $\geq$” follows from the universal property of the absolute filling radius (see the comment after 2.33). To show the inequality $\leq$”, we first recall that
$\rho(\omega;\mathcal{Z})=\rho(\kappa^{\prime}_{*}(\omega);\kappa^{\prime}(\mathcal{Z})\subset L^{\infty}(\mathcal{Z})),$
where $\kappa^{\prime}:\mathcal{Z}\to L^{\infty}(\mathcal{Z})$ is the Kuratowski map. Second, we make use of the injectivity (2.26) of $L^{\infty}(\mathcal{X})$ in the following way. The inclusion $\mathcal{Z}\subset L^{\infty}(\mathcal{X})$, via the Kuratowski identification $\mathcal{Z}\overset{\sim}{\to}\kappa^{\prime}(\mathcal{Z})$, gives rise to a distance preserving (hence, $1$-Lipschitz) map $\kappa^{\prime}(\mathcal{Z})\to L^{\infty}(\mathcal{X})$. This map can be extended to a $1$-Lipschitz map $p:L^{\infty}(\mathcal{Z})\to L^{\infty}(\mathcal{X})$, by invoking the injectivity of $L^{\infty}(\mathcal{X})$. This map pushes forward the cycle $\kappa^{\prime}_{*}(\omega)$ to $\omega$. Moreover, it pushes forward any chain filling $\kappa^{\prime}_{*}(\omega)$ inside $\mathcal{N}_{r}(\kappa^{\prime}(\mathcal{Z}))\subset L^{\infty}(\mathcal{Z})$, to a chain filling $\omega$ inside $\mathcal{N}_{r}(\mathcal{Z}\subset L^{\infty}(\mathcal{X}))$. ∎
To estimate VR-lifespans, we need to know how to upper-bound quantities like $\rho\big{(}\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\big{)}$, for which Theorem 3.3 is not applicable: the set $\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$ is not compact (even after taking the closure). One way to deal with non-compactness is to intersect $\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$ with the closure of the convex hull of $\mathcal{X}$, like we did in 3.11. That theorem, together with the inequality $\operatorname{TW}_{k}(\cdot)\leq\operatorname{AW}_{k-1}(\cdot)$, will give us a bound on $\rho\big{(}\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\big{)}$ in terms of the Alexandrov width of $\overline{\mathcal{N}}_{r}\big{(}\kappa(\mathcal{X})\subset\overline{\mathrm{conv}}(\kappa(\mathcal{X}))\big{)}$. It turns out the latter can be interpreted in terms of the Urysohn width.
@@
In the inequality $\rho(\omega;\mathcal{X})\leq\frac{1}{2}\operatorname{UW}_{k-1}(\mathcal{X})$, the left-hand side is defined via the Kuratowski embedding in $L^{\infty}(\mathcal{X})$, and it is instructive to interpret the right-hand side as well, in a way compatible with the picture of $\kappa(\mathcal{X})$ sitting in $L^{\infty}(\mathcal{X})$. This interpretation goes back to Tikhomirov [Tik76]. Some references in English are [Gro83, Appendix 1, Proposition (D)] by Gromov and [Bal21, Theorem 2.1.9 and Definition 2.3.1] by Balitskiy. Here we provide a short summary. The width $\operatorname{UW}_{k-1}(\mathcal{X})$ can be equivalently defined as the infimal number $\delta>0$ such that there is map $f:\mathcal{X}\to L^{\infty}(\mathcal{X})$ whose image is at most $(k-1)$-dimensional complex in $L^{\infty}(\mathcal{X})$, and such that $\|\kappa(x)-f(x)\|_{\infty}\leq\delta/2$ for all $x\in\mathcal{X}$. In a nutshell, Theorem 3.3 says that one can try killing higher homology of $\mathcal{X}$ by deforming $\kappa(\mathcal{X})$ to a low-dimensional complex inside $L^{\infty}(\mathcal{X})$, and if every point moves by some controlled distance, then we get an estimate for the filling radius. This interpretation can be shortly reformulated as follows:
@@
$\rho\big{(}\omega;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\big{)}\leq\frac{1}{2}\operatorname{UW}_{k-1}\big{(}\overline{\mathcal{N}}_{r}(\kappa(\mathcal{X})\subset\overline{\mathrm{conv}}(\kappa(\mathcal{X})))\big{)}.$
We will go further and estimate $\rho\big{(}\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\big{)}$ even better, by replacing the convex hull with the tight span (2.28). The following result subsumes Theorem 3.3 when $r=0$. The specific realization of $\operatorname{E}(\mathcal{X})$ in $L^{\infty}(\mathcal{X})$ that was described in Section 2.2 has an isometric copy of $\mathcal{X}$ inside, and the corresponding embedding of $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$ agrees with the Kuratowski embedding.
To simplify notation a bit, in the rest of this section, we omit $\kappa$ and simply write $\mathcal{X}\subset\operatorname{E}(\mathcal{X})\subset L^{\infty}(\mathcal{X})$. Notice that the set $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$ is compact, as a closed subset of the compact metric space $\operatorname{E}(\mathcal{X})$.
@@
For any compact metric space $\mathcal{X}$, any integer $k\geq 1$, any positive number $r$, and any homology class $\omega\in\mathrm{H}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))$,
$\rho\big{(}\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\big{)}\leq\frac{1}{2}\operatorname{UW}_{k-1}\big{(}\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}.$
@@
We have a chain of distance-preserving embeddings $\mathcal{X}\subset\operatorname{E}(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ as well as $\mathcal{X}\subset\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. All distances below are measured using the $L^{\infty}$-norm. We need to bound $\rho\big{(}\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)}$.
Recall the following property of the tight span (2.30): there exists a $1$-Lipschitz retraction $\pi:L^{\infty}(\mathcal{X})\to\operatorname{E}(\mathcal{X})$. Our first observation is that $\pi$ maps $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ to $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$. Indeed, take any point $y\in\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Since $\|y-x\|<r$ for some $x\in\mathcal{X}$, and $\pi$ is $1$-Lipschitz retraction, it follows that $\|\pi(y)-x\|=\|\pi(y)-\pi(x)\|\leq\|y-x\|<r$. Therefore, $\pi\big{(}\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\big{)}=\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$. Denote the reverse inclusion
$\iota:\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\hookrightarrow\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})).$
Since both $y$ and $\pi(y)$ lie in the open radius $r$ ball centered at $x$, the whole straight line segment between them lies in that ball and in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Therefore, a cycle representing $\omega$ can be continuously deformed to a cycle representing $\iota_{*}\circ\pi_{*}(\omega)$ by letting each $y$ in the support of $\omega$ slide along the straight line segment towards $\pi(y)$. This homotopy takes place entirely in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$, and therefore, $\omega=\iota_{*}\circ\pi_{*}(\omega)$.
@@
$\rho\big{(}\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)}\leq\rho\big{(}\pi_{*}(\omega);\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)}.$
Indeed, any cycle $S$ representing $\pi_{*}(\omega)\in\widetilde{\mathrm{H}}_{k}\big{(}\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}$ lies entirely in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ and represents there the homology class $\iota_{*}\circ\pi_{*}(\omega)=\omega\in\widetilde{\mathrm{H}}_{k}\big{(}\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\big{)}$. If $S$ bounds in the neighborhood of $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$ of certain radius, then it bounds even in a smaller neighborhood of $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$.
Using the observation at the beginning of this subsection, we deduce that
$\rho\big{(}\pi_{*}(\omega);\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)}=\rho\big{(}\pi_{*}(\omega);\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}.$
The next step follows trivially from the definition of the filling radius:
$\rho\big{(}\pi_{*}(\omega);\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}=\rho\big{(}\pi_{*}(\omega);\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}.$
@@
$\rho\big{(}\pi_{*}(\omega);\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}\leq\frac{1}{2}\operatorname{UW}_{k-1}\big{(}\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}.$
@@
$d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho\big{(}\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)},$
@@
$\lim_{r\to b_{\omega}+0}\operatorname{UW}_{k-1}\big{(}\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}=\operatorname{UW}_{k-1}\big{(}\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}.$